# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 6 Randomized Low-Rank Tensor Decompositions
As in the matrix setting, we are interested in algorithms that compute low-rank tensor decompositions accurately and efficiently, with provable performance guarantees that are substantiated by empirical evidence. Because many tensor algorithms fundamentally rely on classical matrix algorithms, we can frequently fall back on the ideas introduced in Sect. 4 to leverage randomization for efficient tensor decompositions. However, given the rapidly increasing number of applications that involve large-scale tensorial data, it is often impractical or infeasible to rely on basic implementations of the randomized rangefinder, e.g., with Gaussian sketching. The curse of dimensionality for $d$-way arrays is even more pernicious than for matrices, so great care must be taken to develop randomized algorithms that are not only accurate, but also scalable.
In this section, we survey new developments in fast randomized algorithms for low-rank tensor decompositions in the CP and Tucker formats. We draw connections and refer back to our presented matrix algorithms when relevant, augmenting them with emerging ideas from the recent proliferation of work on randomized algorithms for tensors. Sect. 6.1 explores how CP-ALS can be accelerated with randomization, specifically through fast sampling methods that exploit the structure of the overdetermined least squares problems solved in each iteration. Sect. 6.2 summarizes randomized methods to compute Tucker decompositions having orthogonal factor matrices, which naturally extend the randomized rangefinder to the tensor setting via the standard randomized (ST-)HOSVD and randomized HOOI, as well as a newly proposed algorithm for more efficient sketching. These randomized algorithms are contrasted with those for tensor ID/CUR decompositions in Sect. 6.3, which includes both the CP and Tucker formats. Our intent throughout this section is not to provide an exhaustive summary of all recent developments, but rather to broadly survey them to direct the interested reader to the relevant literature, and then hone in on key ideas that may stimulate future work in these emergent areas.
@@ 6.1 Randomized Algorithms for CP Decomposition
Many randomized algorithms for CP decompositions seek to exploit the structure of the Khatri-Rao product (KRP) in CP-ALS, cf. (47), for fast sketching and sampling. In Wang2015, the TensorSketch algorithm of Pham2013 is used to sketch the input tensor without explicitly forming it, then the KRP is efficiently computed. The “randomized ALS” algorithm of Reynolds2016 incorporates projections onto random tensors to improve the conditioning of the least squares problems in CP-ALS. A procedure to sample rows of the KRP near-optimally via fast leverage score approximations is presented in Cheng2016; however, the least squares problem is not sketched in this work (only the KRP). In BattaglinoBallardKolda2018, a Kronecker fast Johnson-Lindenstrauss transform (KFJLT) is applied to the KRP to reduce the coherence for better sampling performance, cf. Jin2021; MALIK2020120, which is then utilized in a sketched ALS problem without explicitly forming the KRP. The algorithm of Aggour2020 builds on BattaglinoBallardKolda2018 with options for adaptive sketch sizes and regularization terms. We now expand on the recent work of Larsen2022 on approximate leverage score sampling that exploits the structure of the KRP.
@@
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\widehat{\bm{\mathsf{A}}}\bm{\mathsf{Z}}^{\top}-\bm{\mathsf{X}}_{(j)}\right\|^{2}_{\text{F}}=\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{Z}}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\right\|^{2}_{\text{F}},$ (51)
@@
In Larsen2022, a randomized sampling strategy based on fast leverage score approximation, CP-ARLS-LEV, is used to accelerate each least squares sub-problem (51). Namely, $S\ll N^{(-j)}$ rows of $\bm{\mathsf{Z}}$ are sampled according to an approximate leverage score distribution. The corresponding $S$ rows of $\bm{\mathsf{X}}_{(j)}^{\top}$ are then selected for a smaller “sketched” problem whose solution is a good approximation of the solution of (51) with high probability. Sketching reduces the cost of solving the $j$th least squares problem to $O(R^{2}S+RSN_{j})$. However, forming $\bm{\mathsf{Z}}$ or $\bm{\mathsf{X}}$ and computing leverage scores would incur a cost of $O(R^{2}N^{(-j)})$. To avoid this, in CP-ARLS-LEV, the leverage scores are estimated as the product of leverage scores of the respective factor matrices in the KRP, and the KRP structure is exploited so that only the $S$ selected rows of $\bm{\mathsf{Z}}$ are formed explicitly.
To this end, let $j\in[d]$ be fixed. We first observe that there is a bijection mapping each row $i\in[N]$ of $\bm{\mathsf{Z}}$ to a (d1)(d-1)-tuple $(i_{d},\ldots,i_{j+1},i_{j-1},\ldots,i_{1})$, with $i_{k}\in[N_{k}]$, $k\neq j,$ which indexes rows of factor matrices in the KRP:
$\bm{\mathsf{Z}}(i,:)=\bm{\mathsf{A}}^{(d)}(i_{d},:)*\cdots*\bm{\mathsf{A}}^{(j+1)}(i_{j+1},:)*\bm{\mathsf{A}}^{(j-1)}(i_{j-1},:)*\cdots*\bm{\mathsf{A}}^{(1)}(i_{1},:),$ (52)
where * again denotes the Hadamard product. Using (Cheng2016, Theorem 3.3), the leverage scores of the rows of $\bm{\mathsf{Z}}$ are bounded above by the product of the leverage scores of the rows of its constituent factor matrices:
@@
From (53), we define a probability distribution on the rows of $\bm{\mathsf{Z}}$; namely, row $i$ is selected with probability
@@
Computing the probabilities in (54) explicitly would require the Kronecker product of leverage scores, costing $O(N^{(-j)})$ operations. However, (Larsen2022, Lemma 9) establishes that each $i_{k}$ in the (d1)(d-1)-tuple corresponding to row $i$ of $\bm{\mathsf{Z}}$ can be sampled independently, using the leverage scores of the rows of $\bm{\mathsf{A}}^{(k)}$, $k\neq j$, which cost $O(N_{k}R^{2})$ to compute. This gives a multi-index corresponding to the $i$-th row of $\bm{\mathsf{Z}}$ without computing the Kronecker product of leverage scores, and $\bm{\mathsf{Z}}(i,:)$ can be formed in O(R(d1))O(R(d-1)) operations using (52). There are additional modifications that can be made to the CP-ARLS-LEV algorithm to further improve performance; we refer the reader to Larsen2022 for details. The basic pseudocode for this procedure is given in Algorithm 8.
Algorithm 8 Alternating Randomized Least Squares for CP with Leverage Scores (CP-ARLS-LEV)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, number of rank-one components $R$, number of samples $S$
4:for $j=1,\ldots,d$ do
5:  Initialize $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$, the vector of row leverage scores for $\bm{\mathsf{A}}^{(j)}$
6:end for
7:while not converged do
8:  for $j=1,\ldots,d$ do
9:   $I=\emptyset$
10:   while $|I|<S$ do
11:     for $k=1,\ldots,j-1,j+1,\ldots,d$ do
12:      Sample $i_{k}\sim\bm{\mathsf{p}}_{k}$ from $N_{k}]$
13:     end for
14:     Add $i=(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$ to II
15:   end while
16:   Form submatrix $\widehat{\bm{\mathsf{Z}}}=\bm{\mathsf{Z}}(I,:)$
17:   Form submatrix $\widehat{\bm{\mathsf{X}}}_{(j)}^{*}=\bm{\mathsf{X}}_{(j)}^{*}(I,:)$
18:   Compute $\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\|\widehat{\bm{\mathsf{Z}}}\widehat{\bm{\mathsf{A}}}^{*}-\widehat{\bm{\mathsf{X}}}_{(j)}^{*}\|^{2}_{\text{F}}$
19:   Update $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$
20:  end for
21:end while
@@
The CP-ARLS-LEV algorithm enjoys strong theoretical guarantees. Namely, the resulting rank-$R$ CP decomposition is an $\varepsilon$-accurate approximation to the input tensor with high probability when the number of samples $S=O\left(R^{d-1}/\varepsilon\right)$. In particular, CP-ARLS-LEV improves on the sampling complexity required by the KFJLT-based method of Jin2021. A distributed-memory implementation of CP-ARLS-LEV is also developed in Bharadwaj2024, along with a distributed-memory version of an algorithm that performs random walks on a binary tree for fast leverage score estimation instead.
@@
The recent work of Malik2022 improves on the sampling complexity of CP-ARLS-LEV by avoiding exponential dependence on $d$, using the fast leverage score estimation technique of DrineasMagdonMahoneyWoodruff12, but replacing the FJLT with recursive sketching Ahle2020. This strategy yields a sampling distribution closer to the exact leverage score distribution, but performs similarly to CP-ARLS-LEV in practice, and its implementation details are beyond the scope of the current manuscript; we refer the reader to Ahle2020; Malik2022. There has also been success in using stochastic gradient descent on the least squares objective function in (51), cf. KoldaHong2020; Wang2023; Yu2024. Randomized algorithms to compute low-rank Tucker decompositions as a pre-processing step for CP-ALS have also been investigated Erichson_2020; Yang2018; Zhou2014. We discuss several of these algorithms in the next section.
6.2 Randomized Algorithms for Tucker Decompositions with Orthogonal Factors
Because the Tucker decomposition provides a natural framework for data compression, methods based on the HOSVD (Algorithm 6) and HOOI (Algorithm 7) have been utilized in many data science and scientific computing applications. However, the Tucker decomposition suffers from the curse of dimensionality, thus classical algorithms can be prohibitively expensive for large problems. As randomized algorithms, particularly the randomized SVD (Algorithm 3), have become mainstays for fast and reliable matrix compression, it is natural to consider how randomization can be leveraged for fast and reliable tensor compression in the Tucker format. In this section, we survey recent work on randomized algorithms for Tucker decompositions with orthogonal factor matrices (as opposed to interpolating, cf. Sect. 5.3). We begin with an overview of major developments in randomized Tucker decompositions, before detailing two algorithms based on the randomized SVD: the randomized HOSVD and the randomized HOOI.
@@ 6.2 Randomized Algorithms for Tucker Decompositions with Orthogonal Factors
Randomized algorithms for Tucker decompositions have been investigated in many problem settings, generally involving either randomized sampling or sketching. One of the earliest works on randomized HOSVD and HOOI algorithms is Tsourakakis2010, called the MACH-HOSVD and MACH-HOOI, so named for the work of Achlioptas-McSherry Achlioptas2001 that inspired these randomized sampling algorithms. In MACH-HOSVD and MACH-HOOI, the input tensor is sparsified according to a coin flip for each nonzero entry; with probability $p$, a nonzero entry is kept and reweighted by $1/p$, and with probability 1p1-p, the entry is set to 0. The HOSVD and HOOI algorithms are then performed on the sparsified tensor, outputting a Tucker decomposition that approximates the original tensor well in practice, even for small values of $p$, e.g., $p=0.1$. Leverage score sampling of matrix unfoldings to sparsify the tensor is explored in Perros2015, and Tucker decompositions of sub-sampled tensors are also investigated in Oseledets2008; Vervliet2014BreakingTC. A detailed survey of these methods can be found in AhmadiAsl2021Randomized.
@@
As the randomized rangefinder gained traction in the numerical linear algebra community, many investigations began to explore randomized sketching for fast Tucker decompositions. In Zhou2014, the RandTucker algorithm is presented, which relies on the randomized rangefinder to compute an HOSVD from Gaussian sketches of mode unfoldings. Formal analysis of RandTucker is performed in Erichson_2020, which incorporates power iteration into RandTucker for faster spectral decay; it is noted, though, that uniform sampling instead of Gaussian sketching performs similarly and may be preferable for large problems. Power iteration and randomized sketching are utilized in the related work of CheWeiYan2020; CheWeiYan2021 for randomized HOSVD and HOOI, which build on the rank-adaptive algorithm of CheWei2019 that iteratively computes an orthonormal basis for the range of each mode unfolding with $\varepsilon$-accuracy.
@@
Rank-adaptive versions of the randomized (ST-)HOSVD are also presented in MinsterSaibabaKilmer2020, which apply the adaptive randomized rangefinder to each mode unfolding. In sun2020low, a single-pass algorithm based on the HOSVD is developed that significantly improves on the cost of storing the random sketching matrices, via Khatri-Rao products of random dimension reduction maps (called tensor random projections); details can also be found in TroppStreaming2019. The Sketch-STHOSVD algorithm of Dong2023Sketching is inspired by sun2020low but incorporates power iteration, whereas the Tucker-TS and Tucker-TTMTS algorithms of Malik2018 are single-pass variants of HOOI using TensorSketch pagh2013compressed. Cost-effective sketching for Tucker decompositions is also investigated in BucciRobol2024MultilinearNystrom, which replaces the randomized SVD step in computing approximate bases with generalized Nyström. In the recent work of Hashemi2025, the Randomized Tucker with Single-Mode Sketching (RTSMS) algorithm is introduced as a rank-adaptive method that sequentially builds approximate bases of unfoldings by applying small random sketches one mode at a time, similar to the randomized ST-HOSVD of MinsterSaibabaKilmer2020, which we describe next.
@@
In the HOSVD (Algorithm 6), we compute factor matrices for a Tucker decomposition as the left singular vectors of each mode unfolding. The randomized HOSVD is a natural extension, where we substitute the SVD of each unfolding with the randomized SVD (Algorithm 3). Pseudocode is given in Algorithm 9, this time for the ST-HOSVD to highlight the slight difference in its computation vs. the HOSVD; the R-HOSVD algorithm of MinsterSaibabaKilmer2020 is precisely the HOSVD algorithm, but using the randomized SVD. Note that this algorithm can be made adaptive by using the adaptive randomized rangefinder with a prescribed error tolerance for the low-rank decomposition of each unfolding.
Algorithm 9 Randomized Sequentially Truncated Higher-Order SVD (R-STHOSVD)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, target ranks $R_{1},\ldots,R_{d}$, oversampling parameter $p$
3:Initialize $\bm{\mathcal{G}}=\bm{\mathcal{X}}$
4:for $j=1,\ldots,d$ do
5:  Draw Gaussian matrix $\bm{\mathsf{\Omega}}^{(j)}\in\mathbb{R}^{N^{(-j)}\times(R_{j}+p)}$
6:  Form sketch $\bm{\mathsf{Y}}^{(j)}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}^{(j)}$
7:  $\widehat{\bm{\mathsf{U}}},\widehat{\bm{\mathsf{\Sigma}}},\widehat{\bm{\mathsf{V}}}]=\texttt{svd}(\bm{\mathsf{Y}}^{(j)},R_{j})$
8:  Set $\bm{\mathsf{A}}^{(j)}=\widehat{\bm{\mathsf{U}}}$
9:  Update $\bm{\mathsf{G}}_{(j)}=\widehat{\bm{\mathsf{\Sigma}}}\widehat{\bm{\mathsf{V}}}^{\top}$
10:end for
If $\widehat{\bm{\mathcal{X}}}$ is the output of Algorithm 9 (or the randomized HOSVD, with Gaussian sketching), it is proven in (MinsterSaibabaKilmer2020, Theorem 3.1, 3.2) that the expected error is
$\mathbb{E}_{\{\bm{\mathsf{\Omega}}^{(j)}\}_{j=1}^{d}}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|\leq\left(d+\frac{\sum_{j=1}^{d}R_{j}}{p-1}\right)^{1/2}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}_{opt}\|.$ (55)
Letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity, the computational cost of R-STHOSVD as in Algorithm 9 is $O(\sum_{j=1}^{d}k^{j}n^{d-j+1}+k^{j}n^{d-j})$. By contrast, R-HOSVD (substituting the SVD with the randomized SVD in Algorithm 6) has computational cost $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$. In other words, the cost is reduced roughly by a factor of $n$ when using randomized (ST-)HOSVD vs. deterministic. Algorithmic performance can also be improved in the R-STHOSVD by processing the largest modes first, and while the error of a given output does depend on processing order, the worst-case expected error does not, cf. (MinsterSaibabaKilmer2020, Sect. 3.3).
@@
Recall from Section 5.3 that higher-order orthogonal iteration (HOOI) is a method of computing Tucker decompositions by iteratively solving least-squares problems for each mode’s factor matrix, which is typically more accurate than the HOSVD but more computationally expensive. Similar to our goal in CP-ALS, for a given input $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, we are seeking a rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}$ that minimizes
@@
where $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$ and $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$. Through an appropriate transformation of the objective function, we can re-write the HOOI procedure in Algorithm 7 as two main steps that are repeated until convergence; cf. Malik2018:
@@
    For $j=1,\ldots,d$, solve $\bm{\mathsf{A}}^{(j)}=\underset{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}{\arg\min}\|\bm{\mathsf{Z}}^{(j)}\bm{\mathsf{G}}_{(j)}^{\top}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\|^{2}$, where
@@
We observe that each of these two steps involve large overdetermined least squares problems, and similar to the strategy in CP-ARLS-LEV, we will exploit the matrix product structures to solve smaller sketched least squares problems efficiently, now with TensorSketch, cf. (Malik2018, Algorithm 2 (TUCKER-TS)).
@@
TensorSketch   We briefly summarize TensorSketch pagh2013compressed; Pham2013; AvronNguyenWoodruff2014; DiaoSongSunWoodruff2018, a specialized version of the CountSketch operator of Charikar2004. A CountSketch operator can be defined as a linear map $\bm{\mathsf{S}}:\mathbb{R}^{I}\rightarrow\mathbb{R}^{J}$ given by $\bm{\mathsf{S}}=\bm{\mathsf{P}}\bm{\mathsf{D}}$, where $\bm{\mathsf{P}}\in\mathbb{R}^{J\times I}$ is a matrix with $\bm{\mathsf{P}}(h(i),i)=1$ and all other entries equal to 0, for a random map $h:[I]\rightarrow[J]$ such that $\mathbb{P}[h(i)=j]=1/J$ for all $i\in[I]$, $j\in[J]$. The matrix $\bm{\mathsf{D}}\in\mathbb{R}^{I\times I}$ is a diagonal matrix with each entry equal to $\pm 1$ with equal probability.

### v2

@@ 6 Randomized Low-Rank Tensor Decompositions
As in the matrix setting, we are interested in algorithms that compute low-rank tensor decompositions accurately and efficiently, with provable performance guarantees, and with demonstrated high practical speed. Because many tensor algorithms fundamentally rely on classical matrix algorithms, we can frequently fall back on the ideas introduced in Sect. 4 to leverage randomization for efficient tensor decompositions. However, given the rapidly increasing number of applications that involve large-scale tensorial data, it is often impractical or infeasible to rely on basic implementations of the randomized rangefinder, e.g., with Gaussian sketching.
In this section, we survey new developments in fast randomized algorithms for low-rank tensor decompositions in the CP and Tucker formats. Sect. 6.1 explores how CP-ALS can be accelerated with randomization, specifically through fast sampling methods that exploit the structure of the overdetermined least squares problems. Sect. 6.2 summarizes randomized methods to compute Tucker decompositions with orthogonal factors, which naturally extend the randomized rangefinder to the tensor setting. These algorithms are contrasted with those for tensor ID/CUR decompositions in Sect. 6.3, which includes both the CP and Tucker formats. Our goal is to broadly survey relevant literature and detail recent developments that may stimulate future work.
@@ 6.1 Randomized Algorithms for CP Decomposition
Many randomized algorithms for CP decompositions seek to exploit the structure of the Khatri-Rao product (KRP) in CP-ALS, cf. (41), for fast sketching and sampling. In Wang2015, the input tensor is sketched without being explicitly formed, then the KRP is computed. The “randomized ALS” algorithm of Reynolds2016 incorporates projections onto random tensors to improve the conditioning of the least squares problems in CP-ALS. A procedure to sample rows of the KRP near-optimally via fast leverage score approximations is presented in Cheng2016. In BattaglinoBallardKolda2018, a Kronecker fast Johnson-Lindenstrauss transform (KFJLT, cf. Jin2021) is applied to the KRP to reduce the coherence for better sampling performance. The algorithm of Aggour2020 builds on BattaglinoBallardKolda2018 with options for adaptive sketch sizes and regularization terms. Below, we expand on the recent work of Larsen2022 on improved leverage score sampling that exploits KRP structure.
@@
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\widehat{\bm{\mathsf{A}}}\bm{\mathsf{Z}}^{*}-\bm{\mathsf{X}}_{(j)}\right\|^{2}_{\text{F}}=\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{Z}}\widehat{\bm{\mathsf{A}}}^{*}-\bm{\mathsf{X}}_{(j)}^{*}\right\|^{2}_{\text{F}},$ (45)
@@
In Larsen2022, the CP-ARLS-LEV algorithm is developed to accelerate each least squares sub-problem (45). Namely, $S\ll N^{(-j)}$ rows of $\bm{\mathsf{Z}}$ are sampled according to an approximate leverage score distribution. The corresponding $S$ rows of $\bm{\mathsf{X}}_{(j)}^{*}$ are then selected for a smaller “sketched” problem, whose solution approximates the solution of (45). Sketching reduces the cost of solving the $j$th least squares problem to $O(R^{2}S+RSN_{j})$. However, forming $\bm{\mathsf{Z}}$ or $\bm{\mathsf{X}}$ and computing leverage scores would incur a cost of $O(R^{2}N^{(-j)})$. To avoid this, in CP-ARLS-LEV, the leverage scores are estimated as the product of leverage scores of the respective factor matrices in the KRP, and only the $S$ sampled rows of $\bm{\mathsf{Z}}$ are formed explicitly.
To this end, let $j\in[d]$ be fixed. We first observe that there is a bijection mapping each row $i\in[N^{(-j)}]$ of $\bm{\mathsf{Z}}$ to a (d1)(d-1)-tuple $(i_{d},\ldots,i_{j+1},i_{j-1},\ldots,i_{1})$, with $i_{k}\in[N_{k}]$, $k\neq j,$ which indexes rows of the factor matrices in the KRP:
$\bm{\mathsf{Z}}(i,:)=\bm{\mathsf{A}}^{(d)}(i_{d},:)*\cdots*\bm{\mathsf{A}}^{(j+1)}(i_{j+1},:)*\bm{\mathsf{A}}^{(j-1)}(i_{j-1},:)*\cdots*\bm{\mathsf{A}}^{(1)}(i_{1},:).$ (46)
Using (Cheng2016, Theorem 3.3), the leverage scores of the rows of $\bm{\mathsf{Z}}$ are bounded above by the product of the leverage scores of the rows of its constituent factor matrices:
@@
From (47), we define a probability distribution on the rows of $\bm{\mathsf{Z}}$:
@@
Computing the probabilities in (48) explicitly would cost $O(N^{(-j)})$ operations. However, (Larsen2022, Lemma 9) establishes that each $i_{k}$ in the (d1)(d-1)-tuple can be sampled independently, using the leverage scores of the rows of $\bm{\mathsf{A}}^{(k)}$, $k\neq j$, which cost $O(N_{k}R^{2})$ to compute. This gives a multi-index corresponding to the $i$-th row of $\bm{\mathsf{Z}}$ without computing the Kronecker product of leverage scores, and $\bm{\mathsf{Z}}(i,:)$ can be formed in O(R(d1))O(R(d-1)) operations using (46). There are additional modifications that can be made to CP-ARLS-LEV to further improve performance; see Larsen2022 for details.
@@
The CP-ARLS-LEV algorithm enjoys nice theoretical guarantees. Namely, the resulting rank-$R$ CP decomposition is an $\varepsilon$-accurate approximation to the input tensor when the number of samples $S=O\left(R^{d-1}/\varepsilon\right)$, under certain assumptions on the structure of the randomized DRM (pertaining to its injectivity parameter and the sketched residual norm); cf. (Larsen2022, Appendix A.1). In particular, CP-ARLS-LEV improves on the sampling complexity required by the KFJLT-based method of Jin2021. A distributed-memory implementation of CP-ARLS-LEV is also developed in Bharadwaj2024, along with a distributed-memory version of an algorithm that performs random walks on a binary tree for fast leverage score estimation instead.
@@
The recent work of Malik2022 improves on the sampling complexity of CP-ARLS-LEV by avoiding exponential dependence on $d$, using the fast leverage score estimation technique of DrineasMagdonMahoneyWoodruff12, but using recursive sketching Ahle2020. This strategy yields a sampling distribution closer to the exact leverage score distribution, but performs similarly to CP-ARLS-LEV in practice, and its implementation details are beyond the scope of the current manuscript. There has also been success in using stochastic gradient descent on the least squares objective function in (45), cf. KoldaHong2020; Wang2023; Yu2024. Randomized algorithms to compute low-rank Tucker decompositions as a pre-processing step for CP-ALS have also been investigated; we discuss several of these algorithms next.
6.2 Randomized Algorithms for Tucker Decompositions
The Tucker decomposition provides a natural framework for data compression. As randomized algorithms, particularly the randomized SVD (Algorithm 3), have become mainstays for fast and reliable matrix compression, it is natural to consider how randomization can be leveraged for fast and reliable tensor compression.
@@ 6.2 Randomized Algorithms for Tucker Decompositions
Randomized algorithms for Tucker decompositions have been investigated in many problem settings, generally involving either randomized sampling or sketching. One of the earliest works on randomized HOSVD and HOOI algorithms is Tsourakakis2010, called the MACH-HOSVD and MACH-HOOI. In MACH-HOSVD and MACH-HOOI, the input tensor is sparsified according to a coin flip for each nonzero entry; with probability $p$, a nonzero entry is kept and reweighted by $1/p$, and with probability 1p1-p, the entry is set to 0. Leverage score sampling of matrix unfoldings to sparsify the tensor is explored in Perros2015, and Tucker decompositions of sub-sampled tensors are also investigated in Oseledets2008; Vervliet2014BreakingTC. A detailed survey of these methods can be found in AhmadiAsl2021Randomized.
@@
As the randomized rangefinder gained traction in the numerical linear algebra community, many investigations began to explore randomized sketching for fast Tucker decompositions. In Zhou2014, the RandTucker algorithm is presented, which relies on the randomized rangefinder to compute an HOSVD from Gaussian sketches of mode unfoldings. Formal analysis of RandTucker is performed in Erichson_2020, which incorporates power iteration into RandTucker for faster spectral decay; it is noted, though, that uniform sampling instead of Gaussian sketching performs similarly and may be preferable for large problems. Power iteration and randomized sketching are utilized in the related work of CheWeiYan2020; CheWeiYan2021 for randomized HOSVD and HOOI, which build on the rank-adaptive algorithm of CheWei2019 that computes a basis for the range of each mode unfolding with $\varepsilon$-accuracy.
@@
In sun2020low, a single-pass algorithm based on the HOSVD is developed that significantly improves on the cost of storing the random sketching matrices, via Khatri-Rao products of random dimension reduction maps (called tensor random projections); details can also be found in TroppStreaming2019. The Sketch-STHOSVD algorithm of Dong2023Sketching is inspired by sun2020low but incorporates power iteration, whereas the Tucker-TS and Tucker-TTMTS algorithms of Malik2018 are single-pass variants of HOOI using TensorSketch pagh2013compressed. Cost-effective sketching for Tucker decompositions is also investigated in BucciRobol2024MultilinearNystrom, which replaces the randomized SVD step in computing approximate bases with generalized Nyström. In the recent work of Hashemi2025, the Randomized Tucker with Single-Mode Sketching (RTSMS) algorithm is introduced as a rank-adaptive method that sequentially builds approximate bases of unfoldings by applying small random sketches one mode at a time, similar to the randomized ST-HOSVD of MinsterSaibabaKilmer2020, which we describe next.
@@
The randomized HOSVD is a natural extension of the HOSVD, where we substitute the SVD of each unfolding with the randomized SVD (Algorithm 3). The R-HOSVD algorithm of MinsterSaibabaKilmer2020 is precisely Algorithm 6, but with a sketch instead of the unfolding.
If $\widehat{\bm{\mathcal{X}}}$ is the output of the randomized HOSVD, with Gaussian sketching of each mode-$j$ unfolding, the expected error is within $\approx\sqrt{d(k+1)}$ of the optimal rank-$(k,\ldots,k)$ Tucker approximation error, cf. (MinsterSaibabaKilmer2020, Theorem 3.1). Letting $n$ be the uniform mode size again, the computational cost of R-STHOSVD is $O(\sum_{j=1}^{d}k^{j}n^{d-j+1}+k^{j}n^{d-j})$, where the first term is the cost of computing SVDs of mode unfoldings and the second term is the cost of forming the core tensor. By contrast, the R-HOSVD algorithm costs $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$ operations. In other words, the cost is reduced roughly by a factor of $n$ with Gaussian sketching. Performance can be improved by processing the largest modes first; while the resulting error does depend on processing order, the worst-case expected error does not, cf. (MinsterSaibabaKilmer2020, Sect. 3.3).
@@
Recall from Section 5.3 that HOOI computes Tucker decompositions iteratively by solving least-squares problems in each mode for the factor matrix, generally more accurate than the HOSVD but more expensive. Similar to our goal in CP-ALS, for a given input $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, we are seeking $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}$ that minimizes
@@
where $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$ and $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$. We can re-write the HOOI procedure in Algorithm 7 as two main steps that are repeated until convergence; cf. Malik2018:
@@
    For $j=1,\ldots,d$, solve $\bm{\mathsf{A}}^{(j)}=\underset{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}{\arg\min}\|\bm{\mathsf{Z}}^{(j)}\bm{\mathsf{G}}_{(j)}^{*}\bm{\mathsf{A}}^{*}-\bm{\mathsf{X}}_{(j)}^{*}\|^{2}$, where
@@
We observe that each of these two steps involve large overdetermined least squares problems, and similar to the strategy in CP-ARLS-LEV, we will exploit the matrix product structures to solve smaller sketched least squares problems efficiently, now with TensorSketch.
@@
TensorSketch   TensorSketch pagh2013compressed; Pham2013; AvronNguyenWoodruff2014; DiaoSongSunWoodruff2018 is a specialized version of CountSketch. A CountSketch operator can be defined as a linear map $\bm{\mathsf{S}}:\mathbb{R}^{I}\rightarrow\mathbb{R}^{J}$ given by $\bm{\mathsf{S}}=\bm{\mathsf{P}}\bm{\mathsf{D}}$, where $\bm{\mathsf{P}}\in\mathbb{R}^{J\times I}$ is a matrix with $\bm{\mathsf{P}}(h(i),i)=1$ and all other entries zero, for a random map $h:[I]\rightarrow[J]$ such that $\mathbb{P}[h(i)=j]=1/J$ for all $i\in[I]$, $j\in[J]$. The matrix $\bm{\mathsf{D}}\in\mathbb{R}^{I\times I}$ is a diagonal matrix of $\pm 1$ with equal probability.